Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufactu11. How much manufacturing overhead was applied from the Fabrication Department to Job P and how much was applied to Job Q? (12. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations.) Unit product cost73. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final ans14. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establ15. What was Sweeten Companys cost of goods sold for March? (Do not round intermediate calculations.) Cost of goods sold
Answer #1

Departmental overhead rate :

Molding department = 12750/2500+2.50 = 7.60 per hour

Fabrication department = 16650/1500+3.30 = 14.40 per hour

11) Manufacturing overhead applied

Job P Job Q
Fabrication department 1700*14.40 = 24480 2000*14.40 = 28800

12) Unit product cost

Job P
Direct material 24000
Direct labor 29800
Molding department overhead (2800*7.60) 21280
Fabrication department overhead 24480
Total Cost 99560
Unit 20
Unit product cost 4978

13) Unit product cost

Job Q
Direct material 13500
Direct labor 11900
Molding department overhead (1900*7.60) 14440
Fabrication department overhead 28800
Total Cost 68640
Unit 30
Unit product cost 2288

14) Total price

Job P Job Q
Direct material 24000 13500
Direct labor 29800 11900
Molding department overhead 21280 14440
Fabrication department overhead 24480 28800
Total Cost 99560 68640
Total price 99560*1.8 = 179208 68640*1.8 = 123552
Selling price per unit 8960 4118

15) Cost of goods sold = 99560+68640 = 168200